Cost Behavior

SmokeCity, Inc., manufactures barbeque smokers. Based on past experience, SmokeCity has found that its total annual overhead costs can be represented by the following formula: Overhead cost = $526,300 + $1.49X, where X equals number of smokers. Last year, SmokeCity produced 19,000 smokers. Actual overhead costs for the year were as expected.

Required:

1. What is the driver for the overhead activity?
Number of smokers

For questions 2-4, Enter the final answers rounded to the nearest dollar.

2. What is the total overhead cost incurred by SmokeCity last year?

$

3. What is the total fixed overhead cost incurred by SmokeCity last year?

$

4. What is the total variable overhead cost incurred by SmokeCity last year?

$

For questions 5-7, round your answers to the nearest cent. Use those rounded figures in subsequent computations, if necessary.

5. What is the overhead cost per unit produced?

$ per unit

6. What is the fixed overhead cost per unit?

$ per unit

7. What is the variable overhead cost per unit?

$ per unit

8. Recalculate Requirements 5, 6, and 7 for the following levels of production: (a) 18,600 units and (b) 20,400 units. Round your answers to the nearest cent.

18,600 Units 20,400 Units
Unit cost $ $
Unit fixed cost
Unit variable cost

The reason the unit costs changed in the way they did is because:
Answer #1

Answers

  • #1 CORRECT.
  • #2 Total Overhead = $ 526300 + ($1.49 x 19000 smokers)
    = 526300 + 28310
    = $ 554,610
  • #3
    = $ 526300
  • #4
    Total Variable cost = 1.49 x 19000 = $ 28310
  • #5
    Overhead cost per unit = $ 554610 / 19000 smokers = $ 29.19
  • #6
    Fixed cost per unit = $ 526300 / 19000 = $ 27.7
  • #7
    Variable overhead cost per unit = $ 1.49
  • #8

18,600 Units

20,400 Units

Unit cost

$29.79

$27.29

Unit fixed cost

$28.30

$25.80

Unit variable cost

$1.49

$1.49

Working

18,600 Units

20,400 Units

Unit cost

=28.3+1.49

=25.8+1.49

Unit fixed cost

=526300/18600

=526300/20400

Unit variable cost

1.49

1.49